Prove that the roots of 2x^2+8x+5 are real and irrational
Answer by Theo(13342)

look at the discriminant.
a = 2
b = 8
c = 5
b^2 - 4ac = 64 - 4*2*5 = 64 - 40 = 24
24 is positive and not a perfect square so the roots are real and irrational.
in order for the roots to be real and rational, the discriminant has to be positive and has to be a perfect square.
a perfect square is when the square root of a number is an integer.
sqrt(24) is not an integer.
but it is positive.